Problem 1

Adam Holmes is the Processing Manager of Empire Mortgage Company, a firm that processes loan applications for a number of regional builders. Home buying and therefore mortgage processing is a highly seasonal business, and requires temporary staff during busy processing periods. Holmes hires staff on a monthly basis from two different temporary staffing firms - Professional Temps (PT) and Support on Demand (SD). In June, Empire hired 14 staff members from PT and 10 from SD. PT is a more established firm and SD is a newly organized firm in the staffing market. Holmes has compiled the following information for June:

Budgets for June

PT staff

SD staff

Budgeted hourly rate

$50

$45

Budgeted time per app. (hours)

1.2

1.4

Actual results for June

PT staff

SD staff

Actual hourly rate

$52

$47

Actual time per app. (hours)

1.4

1.2

Number of actual apps completed

2,604

1,600

Questions:

  1. Determine the labor rate and efficiency variances for (a) the 14 PT staff and (b) the 10 SD staff hired in June.

  2. Comment on the efficiency of the PT and SD staff hired by Empire Mortgage

NOTE-- this has been answered by someone on this site but made no sense to me, there is zero explanation just math. Please explain what you are doing. They are saying the numbers than unfavorable but what are you comparing this against?
Answer #1

Labor Rate Variance :Labor Rate Variance is the used to measure difference between the actual cost of direct labor and the standard cost of direct labor to do the actual work during a period

Formula = (Actual rate - Standard rate) x Actual hours worked = Labor rate variance

For PT Staff:

($ 52 - $ 50) * 3645.60 hours = $ 7291.20 Adverse

Note: Working for actual hours worked:

Number of actual apps completed

2604

Actual time per app. = 1.4 hours

Actual hours worked = 2604 * 1.4 hours = 3645.60 hours

For SD Staff:

($ 47 - $ 45) * 1920 hours = $ 3840 Adverse

Note: Working for actual hours worked:

Number of actual apps completed

1600

Actual time per app. = 1.2 hours

Actual hours worked = 1600 * 1.2 hours = 1920 hours

Labor Efficiency Variance: Labor Efficiency Variance is the used to measure difference between the standard cost of actual number of direct labor hours during a period and the standard hours of direct labor for the actual output achieved during the period.

Formula : (Actual hours - Standard hours for actual output) x Standard rate = Labor efficiency variance

For PT Staff:

(3645.60 - 3124.80 ) * $ 50 = $ 26040 Adverse

Note: Calculation of standard hours for actual output:

Actual output = 2604 application

Standard hours to complete one application: 1.2 hours

Hence, standard hours for actual output: 2604 * 1.2 hours = 3124.80 hours

For SD Staff:

(1920 - 2240) * $ 45 = $ 14400 Favorable

Note: Calculation of standard hours for actual output:

Actual output = 1600 application

Standard hours to complete one application: 1.4 hours

Hence, standard hours for actual output: 1600 * 1.4 hours = 2240 hours

Comment:

Since the PT staff has taken more time than what is budgeted hence it is less efficient whereas the SD had completed the job with in less time than what is budgeted hence they are more efficient.